OpenStudy (anonymous):

Sqrt[98 w^12 x^68] simplify the expression assuming x and w may be negative

OpenStudy (anonymous):

7w^6x^34 sqrt(2)

OpenStudy (anonymous):

when your square rooting letters, if their exponents are even just divide by 2, if they are odd divide by 2 but leave the remainder of 1 underneath...... ex: sqrt(x^15) = x^7 sqrt(x^1) ...... ex #2 sqrt(y3h19), y^1h^9 sqrt(y^1h^1)

OpenStudy (anonymous):

the number 98 is a different story brake it down into 2 x 49 we know the square root of 49 is 7 thus that comes on the outside and the 2 stays underneath the squareroot

OpenStudy (anonymous):

\[7 \sqrt{2} w^6 x^{34} \]

OpenStudy (anonymous):

Same as johnnyBreuer .

hero (hero):

The fact that w and x are negative doesn't matter because -1*-1 = 1

hero (hero):

If say, only x was negative, then we would have a problem